[ ESO ]
CRIRES Quality Control:
dark exposures

CAL | HC | refs | QC
Trending & QC1
Data Packages
Data Management
QC links:
dark level
QC1 database (advanced users): browse | plot
   Click on CURRENT to see the current trending (Health Check).
   Click on HISTORY to see the historical evolution of the trending.

Dark frames are measured daily as part of the calibration plan in sets of 3 frames for each DIT value. Darks with DIT = 1 and 300 sec are taken every day as health check. Darks with other DITs are only measured when they are triggered by settings used during the night.

DARK frame. Example frame for detector #3 with DIT = 1 sec. Detector glow is present in the corners close to the amplifiers.


top Dark level and read-out noise

All QC1 parameters for CRIRES dark exposures are measured on the central area with x = 256 ... 768 and y = 128 ... 384 for each detector array separately in order to avoid contamination from the detector glow.

QC1 parameters

parameter QC1 database: table, name procedure
median dark level crires_dark, darkmed The median level of each input raw frame is determined; the QC parameter is the average of the three values.
read-out noise (RON) crires_dark, ron12
crires_dark, ron23
The difference of the second and first (RON1) and of the third and second (RON2) input frame is calculated. The read-out noise is calculated from the difference image as follows:
- generate 100 13x13 pixel windows
- compute the pixel standard deviation in each window
- the read-out noise is the median of these measurements multiplied with sqrt(NDIT/2).
RON2 is the more reliable determination.
standard deviation crires_dark, dark_stddev This is the standard deviation of the median levels of the three input frames.


Dark level and read-out noise depend on the DIT. Both parameters are trended for the daily health check settings with DIT =1 and 300 sec for each detector (chip) separately. For the read-out noise, the ron23 value is chosen.


2010-09: ron on detectors 3 and 4 increased after intervention
2011-04-14: temperature of detectors 3 and 4 adjusted, ron and dark levels back to normal
2011-10-11: ron decreased on detectors 3 and 4 after intervention because of low detector temperatures; fixed on 2011-10-17